# Open X-Embodiment: Robotic Learning Datasets and RT-X Models

### v6

@@ I Introduction
00footnotetext: 1Allen Institute for AI; 2Arizona State University; 3California Institute of Technology; 4Carnegie Mellon University; 5Columbia University; 6EPFL; 7ETH Zürich; 8Flexiv Robotics; 9Georgia Institute of Technology; 10German Aerospace Center; 11Google DeepMind; 12Google Research; 13IO-AI TECH; 14Imperial College London; 15Intrinsic LLC; 16Istituto Italiano di Tecnologia; 17Korea Advanced Institute of Science & Technology; 18Max Planck Institute; 19Meta AI; 20Mila Quebec; 21NVIDIA; 22New York University; 23Princeton University; 24Queensland University of Technology; 25RIKEN; 26Shanghai Jiao Tong University; 27Stanford University; 28Technische Universität Darmstadt; 29The University of Texas at Austin; 30The University of Tokyo; 31Toyota Research Institute; 32University of California, Berkeley; 33University of California, Davis; 34University of California, San Diego; 35University of Edinburgh; 36University of Freiburg; 37University of Illinois Urbana-Champaign; 38University of Michigan; 39University of Montreal; 40University of Pennsylvania; 41University of Southern California; 42University of Technology, Nuremberg; 43University of Texas at Austin; 44University of Washington

### v7

@@ I Introduction
00footnotetext: 1Allen Institute for AI; 2Arizona State University; 3California Institute of Technology; 4Carnegie Mellon University; 5Columbia University; 6EPFL; 7ETH Zürich; 8Flexiv Robotics; 9Georgia Institute of Technology; 10German Aerospace Center; 11Google DeepMind; 12Google Research; 13IO-AI TECH; 14Imperial College London; 15Intrinsic LLC; 16Istituto Italiano di Tecnologia; 17Korea Advanced Institute of Science & Technology; 18Max Planck Institute; 19Meta AI; 20Microsoft Research; 21Mila Quebec; 22NVIDIA; 23New York University; 24Princeton University; 25Queensland University of Technology; 26RIKEN; 27Shanghai Jiao Tong University; 28Stanford University; 29Technische Universität Darmstadt; 30The University of Texas at Austin; 31The University of Tokyo; 32Toyota Research Institute; 33University of California, Berkeley; 34University of California, Davis; 35University of California, San Diego; 36University of Edinburgh; 37University of Freiburg; 38University of Illinois Urbana-Champaign; 39University of Michigan; 40University of Montreal; 41University of Pennsylvania; 42University of Southern California; 43University of Technology, Nuremberg; 44University of Texas at Austin; 45University of Washington